OpenStudy (anonymous):

Joseph has 24 cousins. He knows that he has 6 more female cousins than male cousins. How many female cousins does he have? Solve and show all working out.

OpenStudy (thesecret20111):

Write 2 equations for this Let x = Number of female cousins, y = male cousins equation 1: x=y+6 equation 2: x+y=24 Now solve for x using substitution.

OpenStudy (anonymous):

What do I need to substitute?

OpenStudy (thesecret20111):

Equation 1 tells you the x=y+6.. Substitute this into equation 2 and you get (y+6)+y=24 Now solve that equation for y then we can substitute y into the first equation and solve for x